5. An SPT test was performed at a depth of 4 meters below the ground surface and the following values were obtained: 5-8-10 (each value over a 15 cm depth increment). The efficiency of the SPT hammer was measured as 70%. The unit weight of the soil from 0 to 2 meters is 19.7 kN/m3 , the unit weight of the soil from 2 to 3 meters is 18.6 kN/m3 and the unit weight of the soil from 3 to 4 meters feet is 19.2 kN/m3 . The water table was encountered at a depth of 2.5 meters.

a) Using (N1)60=CN*N60 calculate the N160 value for the sample at 4 meters.

(

b) Assume the SPT sample at 4 meters consisted of a clean, fine-grained sand (SP). Using Equation drained friction angle= ((15.4(N1)60)^.5,)+20

estimate the drained friction angle for the SPT sample
Answer #1

Solution At 8=19.7 kN/m3 - 2 m 12.5m Depth of 4m. value obserbed, (5-8-10) - N = 8+10 ->(N=18] 8 = 18.6 kNm3 W .T . 3m 8 - 19

____________________________________________________________

PS : For further help feel free to contact and please give thumbs up if it helpful to you.